• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ramesh/Kai/Diancheng

    Log entry time 18:29:44 on November 09, 2009

    Entry number 298604

    This entry is a followup to: 298188

    keyword=Access work

    We added a bunch of 16 ns cables by replacing the existing 4 ns ones in
    group 1 to 4 in the section between preshower (actually pion rejector1)
    to total shower (actually the sum of both pion rejectors) and the pion
    rejctor2 to total shower. We also labeled all the newly added cables as
    well as the cables that were added in the same groups the day before
    yesterday (see halog 298188).

    We also checked whether there was any double pulsing in group3 preshower
    Electon Narrow. We saw a double pulsing in a signal coming from medium
    threshold discriminator but it would not be coincident with the signal
    output of the high threshold discriminator so as to form an ANDed output
    to give the Electron Narrow signal. We moved the signal around in other
    available channels in the same discriminator; this did not solve the
    problem of double pulsing. It was not possible to check the double
    pulsing in Elecron Narrow by itself since this signal was too rare with
    cosmic (need to wait several minutes to see a pulse).

    We also checked a pion veto in RHRS that goes to a fastbus TDC. We can
    see it on the scope. We have now plugged in two pion veto signals to the
    TDC module just for a cross-check. These signals will be in both
    histograms R.pv.t_cor[69] and R.pv.t_cor[94].

    We did not observe any double pulsing in T1. We have replaced a 1ns cable
    by a 6ns cable to a T1 signal that goes to the veto circuit so that on an
    average the gas-cer signal could be about 10 ns earlier than T1.